# Jiaming Wei    Dihan Zou

### v1

@@ Appendix B Proof of Proposition 1 and 2
For uniqueness, we following the approach of KangWatt24. Suppose there are two distinct optimal mechanisms $(q_{1},\Pi_{1})$ and $(q_{2},\Pi_{2})$ that differs for some positive-measure range of costs. Then each of the mechanism must satisfy IC and the no-subsidy constraint $qP(q)-cq\geq\Pi$. Consider an alternative mechanism $(\frac{q_{1}+q_{2}}{2},\frac{\Pi_{1}+\Pi_{2}}{2})$. The new mechanism satisfies IC because of linearity. It also satisfies the no-subsidy constraint because $qP(q)$ is strictly concave, we have for each $c$
$\frac{q_{1}+q_{2}}{2}P\left(\frac{q_{1}+q_{2}}{2}\right)-c\frac{q_{1}+q_{2}}{2}>\frac{q_{1}P(q_{1})+q_{2}P(q_{2})}{2}-c\frac{q_{1}+q_{2}}{2}\geq\frac{\Pi_{1}+\Pi_{2}}{2}.$
@@
We also note that unlike KangWatt24, our regulation problem is a convex program with non-majorization constraints:
$q(c)P(q(c))-cq(c)-k\geq\int_{c}^{1}q(x)dx.$
Therefore, we cannot proceed using the novel approach provided in KangWatt24, neither can we adopt the approach of KleinerMajorizationECMA.
@@
So far, we know that if the cost density is nonincreasing, there is a unique cutoff $\hat{c}\in(0,\overline{c})$ for the no-subsidy constraint (9) to bite. We first show that $\tau(p)$ increasing in the firm price $p$ is equivalent to showing the direct unit tax $\tau(p(c)):=g(c)/q(c)$ is increasing in $c$ over $\hat{c},\overline{c})$. We need to show that the firm price $p(c)$ is increasing in $c$. By the definition of firm profit, $p(c)=c+\frac{\Pi(c)+k}{q(c)}=c+\frac{\int_{c}^{\overline{c}}q(x)dx+k}{q(c)}$. We get $\frac{dp(c)}{dc}=-\frac{q^{\prime}(c)[\Pi(c)+k]}{[q(c)]^{2}}>0$, i.e., $p(c)$ is increasing in $c$. In particular, we show that $q^{\prime}(c)<0$ shortly.
@@
$).P(q(c))=c+(1-\alpha)\frac{F(c)}{f(c)}+\frac{\Gamma(\hat{c})+\beta}{f(c)}=\phi(c;\Gamma(\hat{c}$)).
Note that $P(q(c))$ is strictly increasing in $c$ because $f(c)$ is nonincreasing. Because $P(\cdot)$ is strictly decreasing, this implies $)]q(c)=P^{-1}[\phi(c;\Gamma(\hat{c}$))] strictly decreases over $\hat{c},\overline{c})$. The expression in (18) becomes
@@
The first two terms $(1-\alpha)\frac{F(c)}{f(c)}+\frac{\Gamma(\hat{c})+\beta}{f(c)}$ is nondecreasing as $f(c)$ is nonincreasing. What remains is to show $\frac{\int_{c}^{1}P^{-1}[\phi(x;\Gamma(\hat{c}))]\,dx+k}{P^{-1}[\phi(c;\Gamma(\hat{c}))]}$ is decreasing. Denote the demand function to be $D:=P^{-1}$, and define the composite function $H:=D\circ\phi_{\Gamma(\hat{c})}$. It is equivalent to show $\frac{\int_{c}^{1}H(x)\,dx+k}{H(c)}$ is decreasing. First, consider the case $k=0$, then $\frac{\int_{c}^{1}H(x)\,dx}{H(c)}$ is strictly decreasing if $H$ is strictly log-concave. [^27] See Theorem 6 in BagBerg05Logconcave. Next, we show that a set of sufficient conditions for $H$ to be strictly log-concave is that $D$ is strictly log-concave, as well as $f$ being log-concave and nonincreasing. By definition, log-concavity of $H(x)$ requires $\frac{\partial^{2}\log H(x)}{\partial x^{2}}<0$, i.e.,
$\frac{[D\cdot D^{\prime\prime}-(D^{\prime})^{2}](\phi_{\Gamma(\hat{c})}^{\prime})^{2}+D\cdot\overbrace{D^{\prime}}^{<0}\cdot\phi_{\Gamma(\hat{c})}^{\prime\prime}}{D^{2}}<0,$
where $D$ is evaluated at $\phi_{\Gamma(\hat{c})}(c)$, for any $c\in[\hat{c},\overline{c})$. Since $D=P^{-1}$ is strictly log-concave, the term $D\cdot D^{\prime\prime}-(D^{\prime})^{2}]$ is negative. A sufficient condition for the whole expression to be negative is then $\phi_{\Gamma(\hat{c})}(c)^{\prime\prime}\geq 0$, or $\phi_{\Gamma(\hat{c})}(c)$ is convex. We argue that nonincreasing and log-concave $f$ guarantees the convexity of $\phi_{\Gamma(\hat{c})}(c)$ for any fixed $\hat{c}$. Taking the second-order derivative of $\phi_{\Gamma(\hat{c})}(x)$, we get
$\phi_{\Gamma(\hat{c})}^{\prime\prime}(x)=(1-\alpha)\frac{-f^{2}\cdot\overbrace{f^{\prime}}^{<0}+F\cdot[2(f^{\prime})^{2}-f^{\prime\prime}\cdot f]}{f^{3}}+[\Gamma(\hat{c})+\beta]\cdot\frac{2(f^{\prime})^{2}-f^{\prime\prime}\cdot f}{f^{3}}$
Log-concavity of $f$ implies $f\cdot f^{\prime\prime}-(f^{\prime})^{2}\leq 0$, ensuring $\phi_{\Gamma(\hat{c})}^{\prime\prime}(x)\geq 0$. We conclude that $H$ is strictly log-concave.
On the other hand, for $k>0$, differentiating $\frac{\int_{c}^{1}H(x)\,dx+k}{H(c)}$ yields
$1+\frac{-H^{\prime}(c)}{H(c)}\left[\frac{\int_{c}^{1}H(x)\,dx}{H(c)}+\frac{k}{H(c)}\right],$
we show that above expression is negative if $k$ is small. The expression being negative is equivalent to
$(-H^{\prime}(c))\cdot k+(-H^{\prime}(c))\cdot\int_{c}^{1}H(x)\,dx-H^{2}(c)<0.$
$H(c)$ being strictly log-concave implies $\int_{c}^{\overline{c}}H(x)\,dx$ is strictly log-concave, leading to $(-H^{\prime}(c))\cdot\int_{c}^{1}H(x)\,dx-H^{2}(c)<0,\forall c\in[\hat{c},\overline{c})$. Since $H^{\prime}(c)\geq 0$ is finite, there exists $\overline{k}>0$ such that for all $k<\overline{k}$, the above inequality hold. To conclude, the sufficient conditions in Proposition 2 ensure the progressive tax condition (18). This finishes the proof of Proposition 2. ∎
@@
$\lambda_{q}^{\prime}(0)=-P(q(0))f(0)-\gamma(0)[P(0)+qP^{\prime}(0)]<0,$
if $q(0)<q_{LF}(0)$, a contradiction to the requirement that $\lambda_{q}(c)\geq 0$ for all $c$. Therefore, it must be that $q(0)\geq q_{LF}(0)$, i.e., there is no flat segment of $q(c)$ preceding laissez-faire segment. We summarize the constrained optimal demand policy as follows:
@@
Suppose that condition (LF) is violated. Furthermore, assume that $f$ is nonincreasing and log-concave, and that $P^{-1}$ is strictly log-concave. Then there exists some $\overline{k}>0$, for each $k\in[0,\overline{k})$, the optimal regulation implements a demand schedule and a pricing strategy, $(q^{*},p^{*})$, partitioned in at most four segments with respect to cost cutoffs $0\leq c_{L}\leq\hat{c}\leq\overline{c}\leq 1$ and a benchmark price $\hat{p}>0$.

### v2

@@ Appendix B Proof of Proposition 1 and 2
For uniqueness, we follow the approach of KangWatt24.272727We also note that unlike KangWatt24, our regulation problem is a convex program with non-majorization constraints: $q(c)P(q(c))-cq(c)-k\geq\int_{c}^{1}q(x)dx.$ Therefore, we cannot proceed using the novel approach provided in KangWatt24, neither can we adopt the approach of KleinerMajorizationECMA. Suppose there are two distinct optimal mechanisms $(q_{1},\Pi_{1})$ and $(q_{2},\Pi_{2})$ that differ for some positive-measure range of costs. Then each of the mechanism must satisfy IC and the no-subsidy constraint $qP(q)-cq\geq\Pi$. Consider an alternative mechanism $(\frac{q_{1}+q_{2}}{2},\frac{\Pi_{1}+\Pi_{2}}{2})$. The new mechanism satisfies IC because of linearity. It also satisfies the no-subsidy constraint because $qP(q)$ is strictly concave, we have for each $c$
$\frac{q_{1}+q_{2}}{2}P\left(\frac{q_{1}+q_{2}}{2}\right)-c\frac{q_{1}+q_{2}}{2}-k>\frac{q_{1}P(q_{1})+q_{2}P(q_{2})}{2}-c\frac{q_{1}+q_{2}}{2}-k\geq\frac{\Pi_{1}+\Pi_{2}}{2}.$
@@
So far, we know if the cost density is nonincreasing, there is a unique cutoff $\hat{c}\in(0,\overline{c})$ for the no-subsidy constraint (9) to bite. We first show that $\tau(p)$ increasing in the firm price $p$ is equivalent to the direct unit tax $\tau(p(c)):=g(c)/q(c)$ increasing in $c$ over $\hat{c},\overline{c})$, which boils down to verifying $p(c)$ is increasing. By the definition of firm profit, $p(c)=c+\frac{\Pi(c)+k}{q(c)}=c+\frac{\int_{c}^{\overline{c}}q(x)dx+k}{q(c)}$. We get $\frac{dp(c)}{dc}=-\frac{q^{\prime}(c)[\Pi(c)+k]}{[q(c)]^{2}}>0$, i.e., $p(c)$ is increasing in $c$. We show that $q^{\prime}(c)<0$ shortly.
@@
$),P(q(c))=c+(1-\alpha)\frac{F(c)}{f(c)}+\frac{\Gamma(\hat{c})+\beta}{f(c)}=\phi(c;\Gamma(\hat{c}$)),
which is strictly increasing in $c$ because $f(c)$ is nonincreasing, implying $)]q(c)=P^{-1}[\phi(c;\Gamma(\hat{c}$))] strictly decreases over $\hat{c},\overline{c})$. The expression in (18) becomes
@@
The first two terms $(1-\alpha)\frac{F(c)}{f(c)}+\frac{\Gamma(\hat{c})+\beta}{f(c)}$ is nondecreasing as $f(c)$ is nonincreasing. What remains is to show $\frac{\int_{c}^{1}P^{-1}[\phi(x;\Gamma(\hat{c}))]\,dx+k}{P^{-1}[\phi(c;\Gamma(\hat{c}))]}$ is decreasing. Denote the demand function to be $D:=P^{-1}$, and define the composite function $H:=D\circ\phi_{\Gamma(\hat{c})}$. It is equivalent to show $\frac{\int_{c}^{1}H(x)\,dx+k}{H(c)}$ is decreasing. First, consider the case $k=0$, then $\frac{\int_{c}^{1}H(x)\,dx}{H(c)}$ is strictly decreasing if $H$ is strictly log-concave. [^28] See Theorem 6 in BagBerg05Logconcave. Next, we show that a set of sufficient conditions for $H$ to be strictly log-concave is that $D$ is strictly log-concave, as well as $f$ being log-concave and nonincreasing. By Theorem 3 in ZouLogConcave (p. 5), a set of sufficient conditions for the composite function $H=D\circ\phi_{\Gamma(\hat{c})}$ to be strictly log-concave is (i) $D$ is strictly log-concave and decreasing; (ii) $\phi_{\Gamma(\hat{c})}$ is monotone and convex. We only need to show that $f$ being log-concave and decreasing leads to convex $\phi_{\Gamma(\hat{c})}$, which is true by directly verifying the second-order derivative:
$\phi_{\Gamma(\hat{c})}^{\prime\prime}(x)=(1-\alpha)\frac{-f^{\prime}\cdot f^{2}+F\cdot[2(f^{\prime})^{2}-f^{\prime\prime}\cdot f]}{f^{3}}+[\Gamma(\hat{c})+\beta]\cdot\frac{2(f^{\prime})^{2}-f^{\prime\prime}\cdot f}{f^{3}}$
Log-concavity of $f$ implies $f\cdot f^{\prime\prime}-(f^{\prime})^{2}\leq 0$, ensuring $\phi_{\Gamma(\hat{c})}^{\prime\prime}(x)\geq 0$. We conclude that $H$ is strictly log-concave. On the other hand, for $k>0$, differentiating $\frac{\int_{c}^{1}H(x)\,dx+k}{H(c)}$ yields
$1+\frac{-H^{\prime}(c)}{H(c)}\left[\frac{\int_{c}^{1}H(x)\,dx}{H(c)}+\frac{k}{H(c)}\right].$
For small $k$, we show that above expression is negative, which is equivalent to $(-H^{\prime}(c))\cdot k+(-H^{\prime}(c))\cdot\int_{c}^{1}H(x)\,dx-H^{2}(c)<0$. $H(c)$ being strictly log-concave implies $\int_{c}^{\overline{c}}H(x)\,dx$ is strictly log-concave, leading to $(-H^{\prime}(c))\cdot\int_{c}^{1}H(x)\,dx-H^{2}(c)<0,\forall c\in[\hat{c},\overline{c})$. Since $H^{\prime}(c)\geq 0$ is finite, there exists $\overline{k}>0$ such that for all $k<\overline{k}$, the above inequality hold. To conclude, the sufficient conditions in Proposition 2 ensure the progressive tax condition (18). This finishes the proof of Proposition 2. ∎
@@
Suppose that condition (LF) is violated, that $f$ is nonincreasing and log-concave, and that $P^{-1}$ is strictly log-concave. Then there exists some $\overline{k}>0$, for each $k\in[0,\overline{k})$, the optimal regulation implements a demand schedule and a pricing strategy, $(q^{*},p^{*})$, partitioned in at most four segments with respect to cost cutoffs $0\leq c_{L}\leq\hat{c}\leq\overline{c}\leq 1$ and a benchmark price $\hat{p}>0$.